# A Universal Deep-Learning Framework for Materials X-ray Absorption Spectra

### v2

@@ IV.5.2 Win Rates
As shown in Table 1, all $w_{\text{s}}$ are larger than 0.5 ranging from 53.37% in Cu FEFF to 71.19% in Mn FEFF, suggesting a consistent spectrum-wise performance improvement of the Tuned-UniversalXAS model over the ExpertXAS model. The overall trend in $w_{\text{s}}$ is in line with $(\eta_{\text{T}}-\eta_{\text{E}})/\eta_{\text{E}}$. It is worth noting that in Cu FEFF, although $(\eta_{\text{T}}-\eta_{\text{E}})/\eta_{\text{E}}$ yields a negative value of -7.42 based on the median MSE, $w_{\text{s}}=53.37\%$ indicates a slight improvement based on more fine-grained spectrum-wise metric. The advantage of the Tuned-UniversalXAS model is further supported by $w_{\text{e}}$, which shows a more significant performance improvement in the energy-wise metric with values varying from 74.74% in Ti VASP to 100% in V, Mn and Co FEFF. Similar to $w_{\text{s}}$, $w_{\text{e}}=75.89\%$ in Cu FEFF, suggesting a significant energy-wise performance improvement of the Tuned-UniversalXAS model over the ExpertXAS model.
@@ IV.5.3 Energy-resolved Improvements
Figure 11: Energy-resolved performance improvement $\tilde{I}(\Delta E)$ when using the Tuned-UniversalXAS model over the ExpertXAS model. The regions of improvement in Tuned-UniversalXAS predictions are highlighted in green, while the regions of degradation are highlighted in red. The spectrum used for normalization ($\mu_{\text{baseline}}$) is overlaid as a dotted line corresponding to the right axis.
@@
Fig. 11 displays $|\tilde{I}(\Delta E)|$ as bar plots for each element, partitioned by the sign of $\tilde{I}(\Delta E)$. To provide context on the practical significance of the improvement, we overlay $\mu_{\text{baseline}}$ with a dotted line corresponding to the y axis the right of each plot. The overlaid $\mu_{\text{baseline}}$ serves as a reference line to cross-relate the performance improvements to regions of practical significance. For all cases, except Cu FEFF and Ti VASP, the green bars dominate the plot, showcasing that the Tuned-UniversalXAS model consistently outperforms the ExpertXAS model across wide rage of energy grid. This aligns with earlier observations based on $w_{\text{e}}$, where Cu FEFF and Ti VASP have the lowest scores. Furthermode, cross-comparison with the average spectrum $\mu_{\text{baseline}}$ reveals that the improvements are more pronounced around regions of the main edge and post-edge that are of high importance in XANES.
@@ V.6 Computational Performance
We envision that the proposed OmniXAS framework provides a practical solution to tackle the computational barrier in XAS spectral analysis. In average, an XASModel takes only a fraction of second (about 0.16 second) to predict a material XANES spectrum, with most of the time spent on featurization (i.e., the M3GNet Block). On the other hand, performing a XANES spectral simulation using VASP requires high-performance computing facilities and an exemplary calculation on Ti K-edge XANES of anatase TiO2 takes about 11.57 CPU hours on Intel(R) Xeon(R) Gold 6336Y processors. Therefore, deep learning XASModels can lead to a remarkable speed up of over five orders of magnitude in XAS prediction. The huge computational speed up of the OmniXAS framework should be comparable or better than typical graph neural network potential frameworks, because first-principles spectral calculations are more expensive than total energy calculations of the ground state. Such a highly efficient XAS prediction capability allows high-throughput XAS modeling applied to a broad material space and opens the door for real-time XAS analysis.
@@ VI Conclusion
By training a single universal model, we successfully addressed the elemental divide often encountered in machine learning models for materials science, establishing a robust foundation for general-purpose XAS predictions across diverse elements.
@@ Acknowledgements.
This research is supported by the U.S. Department of Energy, Office of Science, Office Basic Energy Sciences, Award Number FWP PS-030. This research used theory and computation resources of the Center for Functional Nanomaterials, which is a U.S. Department of Energy Office of Science User Facility, and the Scientific Data and Computing Center, at Brookhaven National Laboratory (BNL) under Contract No. DE-SC0012704. M. R. C. acknowledges BNL Laboratory Directed Research and Development (LDRD) grant no. 24-004. This research also used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC Awards No. BES-ERCAP-20690, 18006, and 14811. We thank Mark Hybertsen for helpful discussions. We would also like to extend our gratitude to Nina Cao and Pavan Ravindra for their valuable feedback on manuscript.

### v3

@@ IV.5.2 Win Rates
As shown in Table 1, all $w_{\text{s}}$ are larger than 0.5 ranging from 53.37% in Cu FEFF to 71.19% in Mn FEFF, suggesting a consistent spectrum-wise performance improvement of the Tuned-UniversalXAS model over the ExpertXAS model. The overall trend in $w_{\text{s}}$ is in line with $(\eta_{\text{T}}-\eta_{\text{E}})/\eta_{\text{E}}$. It is worth noting that in Cu FEFF, although $(\eta_{\text{T}}-\eta_{\text{E}})/\eta_{\text{E}}$ yields a negative value of -7.42 based on the median MSE, $w_{\text{s}}=53.37\%$ indicates a slight improvement based on more fine-grained spectrum-wise metric. The advantage of the Tuned-UniversalXAS model is further supported by $w_{\text{e}}$, which shows a more significant performance improvement in the energy-wise metric with values varying from 74.47% in Ti VASP to 100% in V, Mn and Co FEFF. Similar to $w_{\text{s}}$, a $w_{\text{e}}$ of 75.89% in Cu FEFF suggests a significant energy-wise performance improvement of the Tuned-UniversalXAS model over the ExpertXAS model.
@@ IV.5.3 Energy-resolved Improvements
Figure 11: Energy-resolved performance improvement $\tilde{I}(\Delta E)$ when using the Tuned-UniversalXAS model over the ExpertXAS model. The regions of improvement in Tuned-UniversalXAS predictions are highlighted in green, while the regions of degradation are highlighted in red. The spectrum used for normalization ($\mu_{\text{baseline}}$) is overlaid as a dotted line corresponding to the right y axis.
@@
Fig. 11 displays $|\tilde{I}(\Delta E)|$ as bar plots for each element, partitioned by the sign of $\tilde{I}(\Delta E)$. To provide context on the practical significance of the improvement, we overlay $\mu_{\text{baseline}}$ (dotted line), which serves as a reference line to cross-relate the performance improvements to regions of practical significance. For all cases, except Cu FEFF and Ti VASP, the green bars dominate $|\tilde{I}(\Delta E)|$, showcasing that the Tuned-UniversalXAS model consistently outperforms the ExpertXAS model across wide rage of energy grid. The result also aligns with earlier observations based on $w_{\text{e}}$, where Cu FEFF and Ti VASP have the lowest scores. Furthermore, cross-comparison with the average spectrum $\mu_{\text{baseline}}$ reveals that the improvements are more pronounced around regions of the main edge and post-edge that are of high importance in XANES.
IV.6 Knowledge Transfer Across Chemical Environments
We now examine how the knowledge transfer varies across different local chemical environments. This analysis provides insight into which structural motifs benefit most from the general spectral trends captured by our cascaded transfer learning approach.
To analyze the variation in knowledge transfer effectiveness across different chemical environments, we examine model performance for specific local descriptors (OS, CN, and OCN) previously discussed in Sec. III.2. For each element, we partition both training and test datasets into subgroups based on descriptor values. The environments with fewer than 10 samples are excluded from our analysis to ensure statistical reliability.
For each specific chemical environment characterized by descriptor value $x$ (e.g., coordination number $\text{CN}=6$), we extend our performance metric approach from Sec. IV.1 by defining a subgroup baseline spectrum. Similar to the baseline spectrum used in the original $\eta$ metric (Eq. 10), this subgroup baseline is calculated as the mean of all training spectra within that specific chemical environment:
$[\mu_{\mathrm{baseline}}]_{j}^{(x)}=\frac{1}{N_{\text{train}}^{(x)}}\sum_{i=1}^{N_{\text{train}}^{(x)}}\mu_{j}^{(i,x)},$ (19)
where $N_{\text{train}}^{(x)}$ is the number of training samples with descriptor value $x$.
For a model $M$ (either ExpertXAS or Tuned-UniversalXAS), we compute the median MSE over test samples in subgroup $x$:
$\xi^{(M,x)}=\mathrm{Median}(\{J^{(i,x)}\}_{i=1}^{N_{\text{test}}^{(x)}}),$ (20)
where $J^{(i,x)}$ is the MSE of the $i$-th spectrum in the test set having descriptor value $x$ out of a total of $N_{\text{test}}^{(x)}$ spectra.
We then define the environment-specific performance metric:
$\eta^{(M,x)}=\frac{\xi_{\mathrm{baseline}}^{(x)}}{\xi^{(M,x)}},$ (21)
where $\xi_{\mathrm{baseline}}^{(x)}$ is the median MSE of the subgroup baseline model.
To quantify the improvement of Tuned-UniversalXAS ($T$) over ExpertXAS ($E$) for a specific chemical environment characterized by value $x$, we calculate
$\tilde{\eta}^{(x)}=\frac{\eta^{(T,x)}-\eta^{(E,x)}}{\eta^{(E,x)}}\times 100.$ (22)
A positive $\tilde{\eta}^{(x)}$ indicates that Tuned-UniversalXAS outperforms ExpertXAS for the chemical environment characterized by $x$, with larger values representing greater improvement. Conversely, negative values indicate performance degradation where the ExpertXAS model performs better than the Tuned-UniversalXAS model.
Figs. 12 - 14 display $\tilde{\eta}^{(x)}$ across OS, CN and OCN. We can see that the majority of the results in OS (Fig. 12) show an improvements, with the magnitude varying dramatically across elements and its chemical environments. Mn with OS=4 shows exceptional improvement exceeding 250%. Very high improvement near 100% are also found in Mn and Co with OS=3. In other systems, most of the improvements are below 20%. Only a few systems show small negative $\tilde{\eta}^{x}$, e.g., in Mn (OS=2), Fe (OS=2, 4) and Cu (OS=2, 3), where transfer-learning from the universalXAS model slightly adversely affected the model performance. Similarly, CN results in Fig. 13 overall show a consistent improvement, with the max $\tilde{\eta}^{x}$ of about 130% in Mn with CN=6. According to the UMAP analysis in Fig 5g (blue color), CN=6 data form a very diffusive distribution nearly spreading across the entire spectral dataset. This suggests that knowledge transfer in CN=6 could be very efficient. Indeed, a strong correlation between positive $\tilde{\eta}^{x}$ values with CN=6 is identified for all the elements, except for Cu where its CN=6 UMAP cluster is not as obvious as other elements. In comparison to OS and CN, $\tilde{\eta}^{x}$ of OCN exhibit pronounced improvement more broadly distributed at almost all OCN values as shown in Fig. 14. This trend is likely attributed to the very diffusive nature of the OCN spectral UMAP pattern in most of the OCN values as shown in Fig. 5h.
Figure 12: Performance improvement across oxidation states (OS). Positive values (blue bars) indicate Tuned-UniversalXAS outperforms ExpertXAS; negative values (red bars) indicate the opposite.
Figure 13: Performance improvement across coordination numbers (CN). Same format as Fig. 12.
Figure 14: Performance improvement across oxygen coordination numbers (OCN). Same format as Fig. 12.
@@ V.6 Computational Performance
We envision that the proposed OmniXAS framework provides a practical solution to tackle the computational barrier in XAS spectral analysis. In average, an XASModel takes only a fraction of second (about 0.16 second) on an Apple M2 Max chip to predict a material XANES spectrum, with most of the time spent on featurization (i.e., the M3GNet Block). On the other hand, performing a XANES spectral simulation using VASP requires high-performance computing facilities and an exemplary calculation on Ti K-edge XANES of anatase TiO2 takes about 11.57 CPU hours on Intel(R) Xeon(R) Gold 6336Y processors. Therefore, deep learning XASModels can lead to a remarkable speed up of over five orders of magnitude in XAS prediction. The huge computational speed up of the OmniXAS framework should be comparable or better than typical graph neural network potential frameworks, because first-principles spectral calculations are more expensive than total energy calculations of the ground state. Such a highly efficient XAS prediction capability allows high-throughput XAS modeling applied to a broad material space and opens the door for real-time XAS analysis.
@@ VI Conclusion
We demonstrate that the concept of universal XAS model can indeed be established for light transition metals, based on common spectral trends associated with the local chemical descriptors in this class of materials. This is an important initial step towards developing a fully universal ML framework that covers most of the periodic table. By training a single universal model, we successfully addressed the elemental divide often encountered in machine learning models for materials science, establishing a robust foundation for general-purpose XAS predictions across diverse elements.
@@ Acknowledgements.
This research is supported by the U.S. Department of Energy, Office of Science, Office Basic Energy Sciences, Award Number FWP PS-030. This research used theory and computation resources of the Center for Functional Nanomaterials, which is a U.S. Department of Energy Office of Science User Facility, and the Scientific Data and Computing Center, at Brookhaven National Laboratory (BNL) under Contract No. DE-SC0012704. M. R. C. acknowledges BNL Laboratory Directed Research and Development (LDRD) grant no. 24-004. This research also used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC Awards No. BES-ERCAP-20690, 18006, and 14811. We thank Mark Hybertsen and Gerald T. Seidler for helpful discussions. We would also like to extend our gratitude to Nina Cao and Pavan Ravindra for their valuable feedback on manuscript.